Solve equation
(x-5)(4x+5)=0
Show me the steps

Answer by Ave(106)

split it into two equations
x-5=0 add 5 both sides
x=5
and
4x+5=0 add -5 both sides
4x=-5 divide by 4
x=-5/4
{-5/4, 5}